OpenStudy (anonymous):

The volume of a prism is found by multiplying the area of the base by the height. A rectangular prism has a volume of 144 cm3 , and a square base with a width of 3 cm. What is the height of the prism, in cm? A. 4 B. 8 C. 48 D. 16

jimthompson5910 (jim_thompson5910):

What is the area of the base

OpenStudy (anonymous):

48

jimthompson5910 (jim_thompson5910):

we have "a square base with a width of 3 cm."

jimthompson5910 (jim_thompson5910):

which looks like this 

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

that area is _____

OpenStudy (anonymous):

12

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

A = s^2 A = 3^2 A = 9

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

That should help you find the height

OpenStudy (anonymous):

so it is d

jimthompson5910 (jim_thompson5910):

good